OpenStudy (anonymous):

A water ride at a local water park has a ride shaped like a cone that acts like a funnel whereby guests swirl around the cone until they drop through its center. There is one ride for adults and a similar, smaller version for children. If the adult ride has a radius of 32 feet and the child ride has a radius of 24 feet, what is the ratio between the volumes of each ride? Answer 4:3 16:9 64:27 12:9

OpenStudy (dumbcow):

32/24 = 4/3 volume ratio = (radius ratio)^3 (4/3)^3 = 64/27

OpenStudy (anonymous):

thank u for explaining it to